A sum of money in 50 peswas pieces and 10 peswas pieces amounts to ₵60. If there
are, in all, 440 coins, write down the number of 50 peswas and 10 peswas coins.

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52858)

.

On coin problems,  see the lessons
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - Three methods for solving standard (typical) coin word problems
    - More complicated coin problems
    - Advanced coin problems
    - Solving coin problems mentally by grouping without using equations
    - Non-typical coin problems
    - Santa Claus helps solving coin problem
    - OVERVIEW of lessons on coin word problems
in this site.

You will find there the lessons for all levels - from introductory to advanced,
and for all methods used - from one equation to two equations and even without equations.

A convenient place to quickly observe these lessons from a  "bird flight height"  (a top view)  is the last lesson in the list.

Read them and become an expert in solution of coin problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.




Answer by greenestamps(13203)



I will assume from the given numbers that ₵1 = 100 peswas.

You should know how to solve a problem like this using the formal algebraic method the other tutor referred you to.

You can also get valuable mental exercise by solving the problem informally, using logical reasoning and simple arithmetic.

If all 440 coins were 10 peswas, the total value would be ₵44. The actual total value is ₵60, which is ₵16 more than ₵44.

Each 50 peswas coin has a value of 40 peswas more than each 10 peswas coin. So exchanging a 10 peswas coin for a 50 peswas coin increases the total value by 40 peswas.

The number of such exchanges that would be required to make up the additional ₵16 is 1600/40 = 40.

So the number of 50 peswas coins is 40; that means there are 440-40 = 400 10 peswas coins.

ANSWER: 40 coins of 50 peswas; 400 coins of 10 peswas.

CHECK: 40(50)+400(10) = 2000+4000 = 6000 (peswas) = ₵60